A web music store offers two versions of a popular song. The size of the standard version is 2.3 megabytes. The size of the high quality version is 4.1 megabytes. Yesterday, the high quality version was downloaded three times as often as the standard version. The total size download for the two versions was 4234 megabytes. How many downloads of the standard version were there?
Answer by stanbon(75887)

A web music store offers two versions of a popular song. The size of the standard version is 2.3 megabytes. The size of the high quality version is 4.1 megabytes. Yesterday, the high quality version was downloaded three times as often as the standard version. The total size download for the two versions was 4234 megabytes. How many downloads of the standard version were there?
--------------------
Equations:
Quantity Eq:: h = 3s
Megabyte Eq:: 4.1h+2.3s = 4234
---------
Substitute for "h" and solve for "s"::
4.1*3s + 2.3s = 4234
(12.3+2.3)s = 4234
s = 4234/(14.6) = 190 (# of standard version downloads)
h = 3s = 300+270 = 570 (# of high qual version downloaded)
------------
Cheers,
Stan H.
---------